OpenStudy (anonymous):

Determine whether the Mean Value Theorem, Applies to the following functions on the given interval [a,b]. (b) If so, find or approximate the point(s) that are guaranteed to exist by the MEan Value Theorem. f(x) = 3sin(2x); [0,pi/4]

OpenStudy (anonymous):

Would I not be able to use the MVT because the domain on the closed interval [0,pi/4] is every number except at 0?

OpenStudy (anonymous):

Making the function not qualify for being continuous according to the rules of continuity?

OpenStudy (phi):

3 sin(2x) is continous

OpenStudy (anonymous):

Oh wait, I forgot that the domain is all values that x can be without making the function undefined. So now that I have that I can use the MVT to get: \[\frac{3\sin(\pi/2)-3\sin(0)}{\pi/4-0}=\frac{3}{\pi/4}=12/\pi \] So I need to proove that f prime at c is 12/pi.

OpenStudy (anonymous):

If the function is continuous over the closed interval and differentiable over the open interval, then the MVT says there *is* a value of \(c\) between \(0\) and \(\dfrac{\pi}{4}\) such that \[f'(c)=\frac{f\left(\dfrac{\pi}{4}\right)-f(0)}{\dfrac{\pi}{4}-0}\] You have to find this \(c\).

OpenStudy (anonymous):

\[f(x)=3\sin2x~~\Rightarrow~~f'(x)=6\cos2x~~\Rightarrow~~f'(c)=6\cos2c\] So you solve for \(c\): \[6\cos2c=\frac{f\left(\dfrac{\pi}{4}\right)-f(0)}{\dfrac{\pi}{4}-0}\]

OpenStudy (anonymous):

I know you could get rid of the 6 to get: \[=>\cos(2c)=2/\pi\] From there I'm not sure. Could I get rid of the 2 to get: \[=>\cos(c)=1/\pi\]

OpenStudy (anonymous):

\[\begin{align*}6\cos2c&=\frac{f\left(\dfrac{\pi}{4}\right)-f(0)}{\dfrac{\pi}{4}-0}\\\\ &=\frac{3\sin\dfrac{2\pi}{4}-3\sin0}{\dfrac{\pi}{4}}\\\\ &=\frac{3-0}{\dfrac{\pi}{4}}\\\\ 6\cos2c&=\frac{12}{\pi}\\\\ \cos2c&=\frac{2}{\pi}\\\\ 2c&=\cos^{-1}\left(\frac{2}{\pi}\right)\\ c&=\frac{1}{2}\cos^{-1}\left(\frac{2}{\pi}\right) \end{align*}\] You'll need a calculator for this one

OpenStudy (anonymous):

How do you go from: \[\cos2c=2/\pi \] to \[=>2c=\arccos(2/\pi)\]

OpenStudy (phi):

the inverse cos "undoes" the cos if you have cos(x) = a and take the inverse cosine of both sides: \[ \cos^{-1}(\cos(x)) = \cos^{-1}(a) \\ x= \cos^{-1}(a) \]

OpenStudy (anonymous):

oh I see, Thanks.